OpenStudy (anonymous):

Daniel wants to save at least 20 dollars by putting pennies in a jar each day. on the first day he puts one penny into the jar. the second day he puts 2 pennies into the same jar. On he nth day he puts nth pennies into the same jar. how may days will it take for him to have at least 20 dollars in the jar?

OpenStudy (zzr0ck3r):

\(\sum_1^ni = \frac{n(n+1)}{2}\) so \(20\) dollars is \(2000\) pennies. so \[\frac{n(n+1)}{2}=2000\] Solve for n

OpenStudy (zzr0ck3r):

I think....

OpenStudy (anonymous):

the question is how many days will it take/